Dog Up! Franks is looking at a new sausage system with an installed cost of $438,000. This cost will be depreciated straight-line to zero over the project’s five-year life, at the end of which the sausage system can be scrapped for $69,000. The sausage system will save the firm $129,000 per year in pretax operating costs, and the system requires an initial investment in net working capital of $29,000. If the tax rate is 35 percent and the discount rate is 9 percent, what is the NPV of this project?
Answer #1

Depreciation = Cost of the Project / Useful life years = $438,000 / 5 = $87,600

Cash flow at year 0 = Cost of Project + Investment in NWC = $438,000 + $29,000 = $467,000

Annual Cash Flow(Year 1 - 5) = [Savings in Cost * (1 - t)] + [Depreciation * t]

= [$129,000 * (1 - 0.35)] + [$87,600 * 0.35]

= $83,850 + $30,660 = $114,510

Additional Cash Flow at year 7 = After-tax Salvage Value + Recovery of NWC

= [$69,000 * (1 - 0.35)] + $29,000 = $44,850 + $29,000 = $73,850

NPV = PV of Cash Inflows - PV of Cash Outflows

= [$114,510 * {1 - (1 + 0.09)-5} / 0.09] + [$73,850 / (1 + 0.09)5] - $467,000

= [$114,510 * 3.8897] + [$73,850 / 1.5386] - $467,000

= $445,403.97 + $47,997.43 - $467,000 = $26,401.40